OpenStudy (anonymous):

Which of the following is the equation of a line perpendicular to the line y=(-3/2)x+4 , passing through the point (3,9)?

OpenStudy (anonymous):

K for this, whenever a line is prependicular to another, the slope is the negative reciprocal. Negative Reciprocal means that the slope, -3/2 will turn into the reciprocal (opposite) -2/3, and the it'll switch signs, from negative to positive, or positive to negative. That means our slope is 2/3. Now we do point-slope formula again y-y1=m(x-x1) m=slope which is 2/3 and (x1,y1)= (3,9) Now we pluck in the stuff in and solve :) y-9=2/3(x-3) [First we distribute what's in parenthesis] y-9=2/3x-2 [Now we add 9 to both sides] y=2/3x+7 -> and that's our answer -Hope that helps :)